Suppose you purchased long-term bonds issued by ABC Corporation. The firm promises to pay an 8% annual coupon until maturity and to pay the par value at maturity, which is 10 years from today. The coupon payment per annum will be $ ____ and it is considered the annuity component of a bond and the maturity value of $ ______ is considered the lump sum component of a bond.
Answer #1

Coupon amount = Par value* Coupon rate

Coupon amount = $1000*8%

Coupon amount = $80

---------------------------------------------------

Maturity value of $1000 will be considered as lumpsum amount of bond.

$1000 is the Par value which will be repaid at the end of the maturity period.

Note: Par value is assumed as $1000.
Answer #2

The annual coupon payment is calculated by multiplying the coupon rate by the par value of the bond. In this case, the coupon rate is 8% and the par value is not given, so we cannot calculate the exact coupon payment.

However, we can calculate the maturity value by using the formula:

Maturity value = Par value + (Coupon payment x Annuity factor)

where the annuity factor is calculated using the formula:

Annuity factor = (1 - (1 + r)^(-n)) / r

where r is the annual interest rate and n is the number of periods (in this case, 10 years).

Assuming a par value of $1,000, we can calculate the annuity factor as:

Annuity factor = (1 - (1 + 0.08)^(-10)) / 0.08 = 6.7101

Therefore, the maturity value would be:

Maturity value = $1,000 + ($80 x 6.7101) = $1,536.81

So the coupon payment per annum is $80, and the maturity value is $1,536.81.


answered by: Hydra Master